2. If lung problems like emphysema occur, how it is affected the blood pH of the patient? Explain in detail. 3. 25.0 mL of 0.212 M NaOH is neutralized by 13.6 mL of an HCl solution. What is the molarity (M) of the HCl solution? State the given and needed quantities, state equalities and conversion factors, also the cancellation.
Answer #1

Balanced chemical equation is:

NaOH + HCl ---> NaCl + H2O

Here:

M(NaOH)=0.212 M

V(NaOH)=25.0 mL

V(HCl)=13.6 mL

According to balanced reaction:

1*number of mol of NaOH =1*number of mol of HCl

1*M(NaOH)*V(NaOH) =1*M(HCl)*V(HCl)

1*0.212*25.0 = 1*M(HCl)*13.6

M(HCl) = 0.3897 M

Answer: 0.390 M